# Scalable architecture for trapped-ion quantum computing using RF traps and dynamic optical potentials

### 08 Jan 2024 v2

@@ I Introduction
Trapped ions have ideal properties to be used as qubits for quantum computing (QC); they feature long coherence times, efficient state preparation and detection techniques, and a high degree of connectivity [1, 2, 3, 4, 5]. A quantum register of 1000s of qubits, or more, can be formed, for example, by utilizing an equally spaced crystal of ions in a linear RF Paul trap. Indeed recent years have seen many experimental attempts to work with increasingly larger trapped-ion registers [6, 7, 8, 9].
@@
Here we propose a scalable architecture for QC based on trapped-ion qubits that maintains the advantages of a long ion-crystal while circumventing its challenges. In our proposal, an arbitrarily long ion-crystal is segmented into cells by means of dynamically operated optical potentials, e.g. optical tweezers. In this way, the crystal’s motional mode structure is modified such that heating rates reflect only the cell size, and not $N$. Furthermore, programmable high-fidelity multi-qubit entangling gates can be implemented independently within each cell simultaneously. In addition, we show that mid-circuit measurements, a cornerstone of quantum error correction (QEC) techniques, as well as other central quantum computational tools, are straightforward to implement in an optically segmented ion-crystal.
@@ II Proposed architecture
The key point enabling scalability is that due to the optical segmentation the cells behave as nearly independent quantum registers. Indeed, we show below that both challenges of scaling-up trapped-ion based QCs, i.e., ion-crystal stability and the performance of logical operations, both scale with the cell size, $C$, and not the total number cells, $S$. Thus the dynamically reconfigurable segmentation removes the fundamental limits on the number of cells and on the number of qubits in the ion-crystal, and accordingly on quantum circuit and simulation size.
@@
Figure 1: Architecture for scalable trapped-ion quantum computing using RF traps and dynamic optical potentials. The ion-crystal (vertical direction) is segmented to cells by dynamically applying optical trapping potentials. Sequential applications of parallel quantum operations and reconfiguration of the optical trapping implements a large-scale quantum circuit on the entire ion-crystal. We highlight a section of the quantum circuit and the sequence used. As shown, optical trapping generates $S$ segments, each containing $C$ computational qubits (purple circles) that are separated by $B_{A}$ consecutive optically confined ‘barrier ions (purple full circles). Additional $B_{B}$ ions in each cell (red circles) are used to support additional segmentation configurations of the ion-crystal, and mid-circuit measurements that enable classical feedback (red lines) for the implementation of e.g. quantum error correction. The horizontal direction shows a typical mode of operation of the QC, namely logical operations, involving multi-qubit and single-qubit logical gates, $U_{s,n}^{\left(A\right)}$, are performed in each cell, $s$, at the $n$th computational step. The optical potentials are then switched to a different segmentation configuration, allowing for large-scale connectivity between cells of the previous configuration, $U_{s,n}^{\left(B\right)}$. An intermediate segmentation configuration, in which all $B_{A}$ and $B_{B}$ ions are optically confined, is used to accommodate for mid-circuit measurement, cooling of the ion-crystal and state re-initialization. Additional computational steps ($n+1$, etc.) can be performed as required, thus ultimately guarantying connectivity between all computational qubits.
@@
A general overview of our scalable architecture, is presented schematically in Fig. 1. In favor of a simple presentation we focus on a section of the entire ion-crystal, and a section of the operations used. We explicitly highlight $C=8$ computational qubits (purple circles) per cell. In practice 10s of qubits may be placed in each cell, and the number of cells is theoretically unlimited (vertical dots), forming a large ion-crystal. The cells are separated by $B_{A}$ barrier ions (purple full circles), with the schematic showing $B_{A}=2$. $B_{B}$ additional ions (red circles) are placed in the center of each cell and are used for segment re-configuration and mid-circuit measurements, with the schematic showing $B_{B}=2$. Thus, the ion-crystal contains $N=\left(C+B_{A}+B_{B}\right)S+B_{A}$ trapped ions.
@@
The figure shows a general mode of operation of the QC, with parallel operations along the vertical axis and sequential operations, ordered from left to right, that form computational steps. Specifically, barrier ions are strongly confined with optical trapping potentials (purple full circles), generating the $S=4$ segments in this example (configuration ‘A’). At the $n$th computational step, local and global driving fields simultaneously implement unitary operations, $U_{s,n}^{(A)}$, with $s=1,2,...,S$. These unitary operators implement components of the overall quantum algorithm, and may involve all $C$ computational qubits (purple circles) and $B_{B}$ auxiliary ions (red circles) which are not optically confined in this configuration.
@@
Next all $B_{A}$ and $B_{B}$ barrier ions are optically confined, enabling mid-circuit measurement of the auxiliary ions as well as recooling of the ion-crystal and preparation of ions to be used as auxiliary ions (detailed below). Classical feedback (red lines), based on the mid-circuit measurement results, may be implemented at this step in order to influence the next layer of unitary evolution.
@@
Next, the segmentation is reconfigured by strongly confining previously unconfined $B_{B}$ barrier ions and removing the confinement of the $B_{A}$ barrier ions, thus interchanging their roles. This dynamically and quickly generates a different configuration of the segments (configuration ‘B’). Unitary operators, $U_{s,n}^{(B)}$, can now be implemented and connect previously uncoupled qubits. Additional measurement and reconfiguration steps can be applied as required. We remark that here we have shown only two basic ion-crystal configurations (A and B), and have further considered here and below $B_{A}=B_{B}$, however various additional segmentation configurations, with different parameters (e.g. segment and barrier sizes), can be flexibly generated provided relevant ions are allocated as barrier ions.
@@ III Application demonstrations
Figure 2: Quantum simulation of a 3D rectangular Ising spin-model on an optically segmented ion-crystal. Computational qubits (purple circles) are coupled in cells (purple brackets) such that each cell holds two planes of the rectangular spin-model, and implements inter-plane (dark-blue and bright-blue links) and intra-plane (red links) couplings (a) Couplings of a single cell, holding two planes of the model, with $d=4$, used for both segmentation configurations. (b) Multiple cells form the 3D Ising model. The two optical configurations (‘A’, ‘B’ and purple brackets) enable interlaced coupling of the planes (solid red and dashed red, respectively), such that two entanglement operations implement the model’s Hamiltonian. (c) We design the control pulses that implement the nearest-neighbor couplings required for the Ising model, here for $d=4$ and $S=5$. The resulting coupling between all qubit pairs (horizontal and vertical axes) is shown (color), exhibiting a block-diagonal structure that implements the links shown in (a), for each cell. (d) Deviation of the qubit-qubit coupling from the ideal structure (log scale), showing a low error, that is not limited by crosstalk between adjacent cells, and enables scaling up. The implementation’s infidelity scales as the deviations squared, and is here evaluated to $10^{-4}$ per cell.
For simplicity, we focus here on a 3D rectangular lattice of spins, though many other geometries can be realized, e.g. hexagonal. Figure 2 shows an implementation of such a $d\times d\times 2S$ rectangular Ising model. Specifically, Fig. 2(a) shows a single cell of the ion-crystal, that contains $2d^{2}$ computational qubits (purple circles), and encodes a pair of two-dimensional planes of the simulated system. Programmable entanglement gates implement the required couplings that generate the model’s geometry, i.e. inter-plane couplings (dark-blue and bright-blue) and intra-plane couplings (red). The model is made three dimensional, as shown in Fig. 2(b), by interlacing the vertical couplings between adjacent layers (solid red and dashed red), using the two segmentation configurations (purple brackets).
As an example, we consider a specific realization of our architecture, namely qubits implemented on the $4S_{\frac{1}{2}}$ Zeeman ground-state manifold of trapped ${}^{40}\text{Ca}^{+}$ ions, coupled by Raman transitions mediated by the $4P$ manifolds. We utilize this realization for a $d=4$ and $S=5$ nearest-neighbor Ising model (see further system details in the SM [47]). We design the control pulses that drive the computational qubits, and generate the required entanglement operations for simultaneous $XX$-coupling of the qubits, in the entire ion-crystal, similarly to the couplings shown in Fig. 2(a). For simplicity, we use a uniform, unity coupling between all nearest-neighbor qubits. Our design protocol is presented and discussed below.
@@
The second example showcases a path towards fault-tolerant quantum computation [48, 49, 50]. As shown in 3(a), each segment (purple brackets, ‘A’) is utilized to encode a single logical error-corrected qubit using a distance-five surface code. The code uses 25 computational qubits to store the logical state, and 12 auxiliary (ancilla) qubits to readout the values of both $X$ and $Z$ stabilizer measurements. The entire set of $X$ stabilizer measurements can be implemented in one step, using a single multi-qubit entanglement gate and mid-circuit measurements. Then, the auxiliary qubits are reinitialized, after which the entire set of $Z$ stabilizers is similarly performed. Clearly cell segmentation does not require 12 barrier ions, thus here most of the 12 auxiliary qubits are implemented by computational qubits in the cell.
Figure 3: Quantum error correction code on an optically segmented ion-crystal. (a) We implement several distance-five logical qubits, shown pictorially as grids of $5\times 5$ qubits (purple circles). Each cell in configuration ‘A’ (purple brackets) houses a single logical qubit (labeled 0 - 12 and 25 - 36) and twelve auxiliary qubits (labeled 13 - 24). Programmable multi-qubit gates generate the required entanglement structure for stabilizer measurements on each logical qubit in parallel. Here we show an implementation that generates $X$-stabilizers (light-green plaquettes and blue lines). $Z$-stabilizers may be implemented with an additional gate (light-red plaquettes). Connectivity across the entire register is enabled by configuration ‘B’ (light-purple brackets) such that each cell in this configuration holds data qubits of two adjacent logical qubits. This enables logical entangling operations between any neighboring encoded qubits via stabilizer measurements formed on the border between the two surface codes (red plaqeuttes and blue lines), using the auxiliary ions in this configuration (labeled 17’-19’). (b) Example of programmable multi-qubit gate generating $X$ and $Z$ stabilizers on five logical qubits, implemented on a segmented ion-crystal with $S=5$, $C=37$ and $B=3$. We design the control pulses that implement the entanglement gates required for the various stabilizers. The resulting qubit-qubit couplings between corresponding qubit pairs (horizontal and vertical axes) is shown (color), exhibiting a block-diagonal structure that reflects the formation of five logical qubits. (c) Deviation of the qubit-qubit coupling from the ideal structure (log scale), showing a low error, that is not limited by crosstalk between adjacent cells, and enables scaling up. The gate’s infidelity scales as the deviations squared, and is here evaluated to $10^{-4}$ per cell.
@@
Entanglement between two logical qubits, implemented on independent surface codes, can be achieved, for example, using ‘lattice surgery’ [51, 52]. In this method, stabilizers formed along the border between two surface codes are measured, projecting the joint state of the logical qubits to an entangled basis. In our case, this is naturally enabled by re-configuring the optical segmentation (light-purple brackets, ‘B’) such that ions from neighboring cells can interact, as shown schematically in Fig. 3 (blue lines and red plaquettes in the center of the ’B’ configuration). Following this operation, the segmentation reverts to its original configuration, and additional stabilizers are measured. The overall result of this procedure is an entangling logical $XX$ operation between the adjacent surface codes.
We emphasize that even though five cells are used in this example, our architecture’s scalability enables using considerably more cells in order to implement many logical qubits. This example shows one potential route to fault tolerance using our architecture; however many different variants could be considered. First the distance of the code, and thus the number of physical qubits required, should be determined by the demands on the overall logical error rate. Moreover, the choice of the surface code itself, while motivated by its high fault tolerance threshold, is not the only possible choice. In fact, there are other high-threshold codes, with potentially better encoding rates, that may be suitable as well [53]. Constructing an optimal protocol for fault tolerant QC within our architecture, including a prescription for performing logical non-Clifford operations, is a subject for further study.
@@ IV Spectral properties of optically-segmented ion-crystals
with $e$ the elementary electron charge, $\epsilon_{0}$ the vacuum permittivity and $d$ is the inter-ion distance of an equidistant ion-crystal. As we show below, a strong optical potential, $\omega_{\text{o.t.p}}>\nu$, drastically changes the motional mode structure of the chain. By incorporating the optical trapping potential with common methods to analyze ion-crystal motion [55], we obtain the normal-mode frequencies and structure in the axial and transverse directions (for further details see the SM [47]).
@@
Transverse modes of motion are used in order to generate entanglement between computational qubits, by mediating qubit-qubit interaction via spin-dependent motion. Typically the entanglement operation duration is inversely related to the frequency difference between adjacent modes of motion [14].
Here, the frequency gap between bands, $\Delta\omega_{b}=\omega_{b}-\omega_{b-1}$, with $\omega_{b}$ the mean frequency of the modes of bulk bands, $b=1,...,C+B_{B}$, marks the typical interaction rate between ions within the same cell. Thus a simple estimation of unwanted crosstalk in the entanglement operations between cells is given by $\varepsilon_{\text{BW},b}=\frac{\text{BW}_{b}}{2\Delta\omega_{b}}$, with $\text{BW}_{b}$ the bandwidth of band $b$, that sets the coupling rate between cells. The resulting estimate is presented in the inset (top left) of Fig. 5 (orange), showing that high $b$ bands, associated with long-wavelength motion across cells, results in a higher crosstalk, at the few percent level. A more precise estimate of the crosstalk is also shown (blue line and green points) and discussed below.
@@
Optically confined ions are not used as computational qubits since optical confinement significantly degrades their coherence due to spontaneous photon scattering and uncontrolled light-shifts. For example, we consider ${}^{40}\text{Ca}^{+}$ ions that are illuminated by a confining optical field at $400$ nm with a beam that has a diameter of 1 $1\text{ }\mu\text{m}$, off-resonant with the $S\leftrightarrow P$ transition. In this setup, a harmonic confinement frequency of $\omega_{\text{o.t.p}}\approx 2.1\nu=(2\pi)*1.75$ MHz, used throughout our analysis, requires an optical power of $20$ mw illuminating each barrier ion. The resulting spontaneous photon scattering rate is approximately $3$ kHz. Photons scattered from barrier ions are significantly detuned (by $10$’s of THz) with respect to the $S\leftrightarrow P$ transition in neighboring computational qubits, which thus acquire only a negligible decoherence rate due to this process.

### 04 Nov 2024 v3

@@ I Introduction
Trapped ions have ideal properties to be used as qubits for quantum computing (QC); they feature long coherence times, efficient state preparation and detection techniques, and a high degree of connectivity [1, 2, 3, 4, 5]. A quantum register of 1000s of qubits, or more, can be formed, for example, by utilizing an equally spaced crystal of ions in a linear RF Paul trap. Indeed recent years have seen many experimental attempts to work with increasingly larger trapped-ion registers [6, 7, 8, 9, 10].
@@
Here we propose a scalable architecture for QC based on trapped-ion qubits that maintains the advantages of a long ion-crystal while circumventing its challenges. In our proposal, an arbitrarily long ion-crystal is segmented into re-configurable cells by means of dynamically operated optical potentials, e.g. optical tweezers. In this way, the crystal’s motional mode structure is modified such that heating rates reflect only the cell size, and not $N$. Furthermore, programmable high-fidelity multi-qubit entangling gates can be implemented independently within each cell simultaneously. In addition, we show that mid-circuit measurements, a cornerstone of quantum error correction (QEC) techniques, as well as other central quantum computational tools, are straightforward to implement in an optically segmented ion-crystal.
@@ II Proposed architecture
The key point enabling scalability is that due to the optical segmentation the cells behave as nearly independent quantum registers. Indeed, we show below that both challenges of scaling-up trapped-ion based QCs, i.e., ion-crystal stability and the performance of logical operations, both scale with the cell size, $C$, and not the total number of cells, $S$. Thus the dynamically reconfigurable segmentation removes the fundamental limits on the number of cells and on the number of qubits in the ion-crystal, and accordingly on quantum circuit and simulation size.
@@
Figure 1: Architecture for scalable trapped-ion quantum computing using RF traps and dynamic optical potentials. The ion-crystal (vertical direction) is segmented into cells by dynamically applying optical trapping potentials. Sequential applications of parallel quantum operations and reconfiguration of the optical trapping implements a large-scale quantum circuit on the entire ion-crystal. We highlight a section of the quantum circuit and the sequence used. As shown, optical trapping generates $S$ segments, each containing $C$ computational qubits (white circles) that are separated by $B_{A}$ consecutive optically confined ‘barrier ions’ (purple filled circles). Additional $B_{B}$ ions in each cell (red circles) are used to support additional segmentation configurations of the ion-crystal, and mid-circuit measurements that enable classical feedback (red lines) for the implementation of e.g. quantum error correction. The horizontal direction shows a typical mode of operation of the QC, namely logical operations, involving multi-qubit and single-qubit logical gates, $U_{s,n}^{\left(A\right)}$, are performed in each cell, $s$, at the $n$th computational step. The optical potentials are then switched to a different segmentation configuration, allowing for large-scale connectivity between cells of the previous configuration, $U_{s,n}^{\left(B\right)}$. An intermediate segmentation configuration, in which all $B_{A}$ and $B_{B}$ ions are optically confined, is used to accommodate for mid-circuit measurement and state re-initialization. Additional computational steps ($n+1$, etc.) can be performed as required, thus ultimately guaranteeing connectivity between all computational qubits.
@@
A general overview of our scalable architecture, is presented schematically in Fig. 1. In favor of a simple presentation we focus on a section of the entire ion-crystal, and a section of the operations used. We explicitly highlight $C=8$ computational qubits (white circles) per cell. In practice 10s of qubits may be placed in each cell, and the number of cells is theoretically unlimited (vertical dots), forming a large ion-crystal. The cells are separated by $B_{A}$ barrier ions (purple filled circles), with the schematic showing $B_{A}=2$. $B_{B}$ additional ions (red circles) are placed in the center of each cell and are used for segment re-configuration and mid-circuit measurements, with the schematic showing $B_{B}=2$. Thus, the ion-crystal contains $N=\left(C+B_{A}+B_{B}\right)S+B_{A}$ trapped ions.
@@
The figure shows a general mode of operation of the QC, with parallel operations along the vertical axis and sequential operations, ordered from left to right, that form computational steps. Specifically, barrier ions are strongly confined with optical trapping potentials (purple filled circles), generating the $S=2$ segments in this example (configuration ‘A’). At the $n$th computational step, local and global driving fields simultaneously implement unitary operations, $U_{s,n}^{(A)}$, with $s=1,2,...,S$. These unitary operators implement components of the overall quantum algorithm, and may involve all $C$ computational qubits (white circles) and $B_{B}$ auxiliary ions (red circles) which are not optically confined in this configuration.
@@
Next all $B_{A}$ and $B_{B}$ barrier ions are optically confined, enabling mid-circuit measurement of the auxiliary ions as well as preparation of ions to be used as auxiliary ions (detailed below). Classical feedback (red lines), based on the mid-circuit measurement results, may be implemented at this step in order to influence the next layer of unitary evolution.
@@
Next, the segmentation is reconfigured by removing the confinement of the $B_{A}$ barrier ions. This dynamically and quickly generates a different configuration of the segments (configuration ‘B’). Unitary operators, $U_{s,n}^{(B)}$, can now be implemented and connect previously uncoupled qubits. Additional measurement and reconfiguration steps can be applied as required. We remark that here we have shown only two basic ion-crystal configurations (A and B), and have further considered here and below $B_{A}=B_{B}$, however various additional segmentation configurations, with different parameters (e.g. segment and barrier sizes), can be flexibly generated provided relevant ions are allocated as barrier ions.
@@ III Application demonstrations
Figure 2: Quantum simulation of a 3D rectangular Ising spin-model on an optically segmented ion-crystal. Computational qubits (white circles) are coupled in cells (purple brackets) such that each cell holds two planes of the rectangular spin-model, and implements inter-plane (dark-blue and bright-blue links) and intra-plane (red links) couplings (a) Couplings of a single cell, holding two planes of the model, with $d=4$, used for both segmentation configurations. (b) Multiple cells form the 3D Ising model. The two optical configurations (‘A’, ‘B’ and purple brackets) enable interlaced coupling of the planes (solid red and dashed red, respectively), such that two entanglement operations implement the model’s Hamiltonian. (c) We design the control pulses that implement the nearest-neighbor couplings required for the Ising model, here for $d=4$ and $S=5$. The resulting coupling between all qubit pairs (horizontal and vertical axes) is shown (color), exhibiting a block-diagonal structure that implements the links shown in (a), for each cell. (d) Deviation of the qubit-qubit coupling from the ideal structure (log scale), showing a low error, that is not limited by crosstalk between adjacent cells, and enables scaling up. The implementation’s infidelity scales as the deviations squared, and is here evaluated to $10^{-4}$ per cell.
For simplicity, we focus here on a 3D rectangular lattice of spins, though many other geometries can be realized, e.g. hexagonal. Figure 2 shows an implementation of such a $d\times d\times 2S$ rectangular Ising model. Specifically, Fig. 2(a) shows a single cell of the ion-crystal, that contains $2d^{2}$ computational qubits (white circles), and encodes a pair of two-dimensional planes of the simulated system. Programmable entanglement gates implement the required couplings that generate the model’s geometry, i.e. inter-plane couplings (dark-blue and bright-blue) and intra-plane couplings (red). The model is made three dimensional, as shown in Fig. 2(b), by interlacing the vertical couplings between adjacent layers (solid red and dashed red), using the two segmentation configurations (purple brackets).
As an example, we consider a specific realization of our architecture, namely qubits implemented on the $4S_{\frac{1}{2}}$ Zeeman ground-state manifold of trapped ${}^{40}\text{Ca}^{+}$ ions, coupled by Raman transitions mediated by the $4P$ manifolds. We utilize this realization for a $4\times 4\times 10$ nearest-neighbor Ising model, i.e. $d=4$ and $S=5$ (see further system details in the SM [48]). We design the control pulses that drive the computational qubits, and generate the required entanglement operations for simultaneous $XX$-coupling of the qubits, in the entire ion-crystal, similarly to the couplings shown in Fig. 2(a). For simplicity, we use a uniform, unity coupling between all nearest-neighbor qubits. Our design protocol is presented and discussed below.
@@
The second example showcases a path towards fault-tolerant quantum computation [49, 50, 51]. As shown in Fig. 3(a), each segment (purple brackets, ‘A’) is utilized to encode a single logical error-corrected qubit using a distance-five surface code. The code uses 25 computational qubits to store the logical state, and 12 auxiliary (ancilla) qubits to readout the values of both $X$ and $Z$ stabilizer measurements. The entire set of $X$ stabilizer measurements can be implemented in one step, using a single multi-qubit entanglement gate and mid-circuit measurements. Then, the auxiliary qubits are reinitialized, after which the entire set of $Z$ stabilizers is similarly performed. Clearly cell segmentation does not require 12 barrier ions, thus here most of the 12 auxiliary qubits are implemented by computational qubits in the cell.
Figure 3: Quantum error correction code on an optically segmented ion-crystal. (a) We implement several distance-five logical qubits, shown pictorially as grids of $5\times 5$ qubits (white circles). Each cell in configuration ‘A’ (purple brackets, ’A’) houses a single logical qubit (labeled 0 - 12 and 25 - 36) and twelve auxiliary qubits (labeled 13 - 24). Programmable multi-qubit gates generate the required entanglement structure for stabilizer measurements on each logical qubit in parallel. Here we show an implementation that generates $X$-stabilizers (light-green plaquettes and blue lines). $Z$-stabilizers may be implemented with an additional gate (light-red plaquettes). Connectivity across the entire register is enabled by configuration ‘B’ (purple brackets, ’B’) such that each cell in this configuration holds data qubits of two adjacent logical qubits. This enables logical entangling operations between any neighboring encoded qubits via stabilizer measurements formed on the border between the two surface codes (red plaquettes and blue lines), using the auxiliary ions in this configuration (labeled 17’-19’). (b) Example of programmable multi-qubit gate generating $X$ and $Z$ stabilizers on five logical qubits, implemented on a segmented ion-crystal with $S=5$, $C=37$ and $B=3$. We design the control pulses that implement the entanglement gates required for the various stabilizers. The resulting qubit-qubit couplings between corresponding qubit pairs (horizontal and vertical axes) is shown (color), exhibiting a block-diagonal structure that reflects the formation of five logical qubits. (c) Deviation of the qubit-qubit coupling from the ideal structure (log scale), showing a low error, that is not limited by crosstalk between adjacent cells, and enables scaling up. The gate’s infidelity scales as the deviations squared, and is here evaluated to $10^{-4}$ per cell.
@@
Entanglement between two logical qubits, implemented on independent surface codes, can be achieved, for example, using ‘lattice surgery’ [52, 53]. In this method, stabilizers formed along the border between two surface codes are measured, projecting the joint state of the logical qubits to an entangled basis. In our case, this is naturally enabled by re-configuring the optical segmentation (purple brackets, ‘B’) such that ions from neighboring cells can interact, as shown schematically in Fig. 3 (blue lines and red plaquettes in the center of the ’B’ configuration). Following this operation, the segmentation reverts to its original configuration, and additional stabilizers are measured. The overall result of this procedure is an entangling logical $XX$ operation between the adjacent surface codes.
We emphasize that even though five cells are used in this example, our architecture’s scalability enables using considerably more cells in order to implement many logical qubits. This example shows one potential route to fault tolerance using our architecture; however many different variants could be considered. First the distance of the code, and thus the number of physical qubits required, should be determined by the demands on the overall logical error rate. Moreover, it is possible to encode two (or more) logical qubits per segment; this would then enable performing logical entanglement operations transversally, as an alternative to lattice surgery. Finally, the choice of the surface code itself, while motivated by its high fault tolerance threshold, is not the only possible choice. In fact, there are other high-threshold codes, with potentially better encoding rates, that may be suitable as well [54]. Constructing an optimal protocol for fault tolerant QC within our architecture, including a prescription for performing logical non-Clifford operations, is a subject for further study.
An in-depth numerical analysis of the QEC circuit can be found in the SM [48]. The analysis focuses on technical aspects of the circuit implementation; it accounts for various heating and cooling mechanisms and provides an estimated resource cost associated with each step in the circuit, and shows that our architecture is experimentally feasible with current state-of-the-art hardware. Furthermore, in our analysis we contrast our architecture with QCCD and photonic-interconnect architectures, showing that our proposal offers up to two orders of magnitude of speed up in implementation of QEC rounds, in terms of overall circuit duration.
@@ IV Spectral properties of optically-segmented ion-crystals
with $e$ the elementary electron charge, $\epsilon_{0}$ the vacuum permittivity and $d$ is the inter-ion distance of an equidistant ion-crystal. As we show below, a strong optical potential, $\omega_{\text{o.t.p}}>\nu$, drastically changes the motional mode structure of the chain. By incorporating the optical trapping potential with common methods to analyze ion-crystal motion [56], we obtain the normal-mode frequencies and structure in the axial and transverse directions (for further details see the SM [48]). In the following, we assume that that the optical trapping potential is equally strong in the axial and a single transverse direction, and zero in the other transverse direction. In general, the relative strength of the potential in each direction can be modified by changing the k-vector of the optical trapping beams.
@@
Transverse modes of motion are used in order to generate entanglement between computational qubits, by mediating qubit-qubit interaction via spin-dependent motion [57]. Typically the entanglement operation duration is inversely related to the frequency difference between adjacent modes of motion [15].
In our case, the frequency gap between bands, $\Delta\omega_{b}=\omega_{b}-\omega_{b-1}$, with $\omega_{b}$ the mean frequency of the modes of bulk bands, $b=1,...,C+B_{B}$, marks the typical interaction rate between ions within the same cell. By contrast, the coupling rate between different cells is determined by the bandwidth, $\text{BW}_{b}$, which is defined as the difference between the highest and lowest frequency modes in band $b$. Thus a simple estimation of unwanted crosstalk between cells during parallel entanglement operations is the ratio of these two rates - given by $\varepsilon_{\text{BW},b}=\frac{\text{BW}_{b}}{2\Delta\omega_{b}}$. The resulting estimate is presented in the inset (top left) of Fig. 5 (orange), showing that high $b$ bands, associated with long-wavelength motion across cells, results in a higher crosstalk, at the few percent level. A more precise estimate of the crosstalk is also shown (blue line and green points) and discussed below.
@@
Optically confined ions are not used as computational qubits since optical confinement significantly degrades their coherence due to spontaneous photon scattering and uncontrolled light-shifts. For example, we consider ${}^{40}\text{Ca}^{+}$ ions that are illuminated by a confining optical field at $400$ nm with a beam that has a diameter of 1 $1\text{ }\mu\text{m}$, off-resonant with the $S\leftrightarrow P$ transition. In this setup, a harmonic confinement frequency of $\omega_{\text{o.t.p}}\approx 2.1\nu=(2\pi)*1.75$ MHz, used throughout our analysis, requires an optical power of $20$ mw illuminating each barrier ion. The resulting spontaneous photon scattering rate is approximately $3$ kHz.
This scattering rate, however, does not pose any fundamental challenges. Photons scattered from barrier ions are significantly detuned (by $10$’s of THz) with respect to the $S\leftrightarrow P$ transition in neighboring computational qubits, which thus acquire only a negligible decoherence rate due to this process. Moreover, because the barrier ions only participate weakly in bulk modes, the heating rate of these modes due to photon scattering is small, $\sim 10^{-2}$ quanta per second (see the SM [48]). It should be noted that the barrier ions themselves do heat as a result of this process; thus, barrier ions are required to be ground state cooled during each reconfiguration step, i.e. before they are released into the bulk. This point is further discussed in section VII.
We remark that it is possible to use different ion species for data qubits and barrier ions. This would allow to easily realize a high optical trapping frequency, by using a closer to resonance trapping field, that scatters many photons, without affecting data qubits. Cooling of such barrier ions is still required, yet becomes simpler as well.